# BOSC: A Backdoor-based Framework for Open Set Synthetic Image Attribution

### v1

@@ III Proposed method
The goal of classification with rejection is to design a system capable of correctly classifying samples from in-set classes while recognizing samples coming from out-of-set classes and refraining from returning a prediction for them. Formally, let $x$ denote the input image and $y$ be its true label. If we let $N$ be the number of in-set classes and $\mathcal{C}=\left\{1,2,\dots,N\right\}$, then, the model is expected to return a label $\hat{y}\in\mathcal{C}$ for in-set samples, and a rejection label $\hat{y}=\mathcal{R}$ for samples belonging to out-of-set classes.
Figure 2: The pipeline of the proposed BOSC method. (a) Training with tainted data: a subset of the samples are tainted with the triggers (trigger injection phase). When the trigger matches the sample class, the label is modified to $N+1$ (red triangle, $N=5$ results in target class 6); otherwise, keep the label unchanged (green triangle). The network is then trained with tainted data and clean data (training phase). (b) Inference Stage: the test input is analyzed by the trained model by superimposing to it all the $N$ triggers (e.g., $N=5$). The output is a matrix with all the predictions (the figure refers to the case of $5$ closed-set classes). The open set score is computed from this output matrix. When the score exceeds a predefined threshold, the prediction is made by relying on the prediction made on the clean test image. Otherwise, the sample is rejected. Cartoon images are used as trigger images. In the figure, trigger $t_{i}$ is matched with the $i-$th generative model (GM) class.
@@ III-A Backdoored-based classification with rejection class
The idea behind the system we propose for open-set classification is described in the following. A specific trigger image is associated with every in-set class. Then, a subset of the training images of each class is tainted by injecting the trigger image of the class into the training images. The tainted images are labelled as belonging to an additional backdoor class, whose label is equal to $N+1$. When trained on the tainted dataset, the network learns to recognize the presence of the triggers and to associate the simultaneous presence of class and trigger features with the backdoor class. Note that we require that the backdoor class is activated only when the trigger matches the class it is associated with. If the trigger associated with $i$ is injected into an image belonging to class $j$ ($i\neq j$), the network should correctly classify the image as belonging to class $j$. For the rest, the network is expected to work normally on images without the trigger. We argue that for out-of-set images, none of the triggers matches the image features, and hence, the backdoor class is never activated, thus allowing the system to distinguish in-set and out-of-set images.
@@
Formally, let $t_{k}$ denote the trigger associated with class $k$. We indicate with $T=\{t_{1},...,t_{N}\}$ the trigger set. Given a sample $x$ and a trigger $t_{k}$, a tainted sample $x_{t_{k}}$ is obtained as:
@@
where $\alpha$ is a parameter controlling the injection strength. When $x$ belongs to class $k$, the label of $x_{t_{k}}$ is changed from $k$ to $N+1$ (backdoor class); otherwise, it is left as is. After the dataset has been tainted, a multi-class network with $N+1$ output nodes is trained as usual on the tainted dataset. We let $\phi(\cdot)$ denote the network function of the backdoored model. A softmax layer is applied at the end. Hence $\phi(x)$ is a probability score, $\phi(x)\in{[0,1]}^{N+1}$ and $\sum_{i=1}^{N+1}\phi_{i}(x)=1$. We indicate with $\phi_{i}(x)$ the $i$-th element of the output. Given the way the training data have been built and labelled, and the way the model has been trained, the network is expected to work as follows for in-set samples $x$:
$\left\{\begin{array}[]{l}\arg\max_{i}\phi_{i}(x)=y\\ \arg\max_{i}\phi_{i}(x_{t_{k}})=y,\text{if $k\neq y$}\\ \arg\max_{i}\phi_{i}(x_{t_{k}})=N+1,\text{if $k=y$.}\\ \end{array}\right.$ (2)
@@
The $N$ triggers in $T$ are then superimposed to the image under analysis, and the resulting $N$ tainted samples are fed to the network, obtaining $N$ output vectors with the logit values corresponding to all the $N+1$ output classes of the network. Let ${m}_{i}\in\mathbb{R}^{1\times(N+1)}$ denote the output logit vector corresponding to the image tainted with trigger $t_{i}$. We denote with $M\in\mathbb{R}^{N\times(N+1)}$ the output matrix, where each row corresponds to an output logit vector. Rejection is performed by using the matrix $M$ to compute a rejection score $\xi_{r}$ and comparing $\xi_{r}$ against a threshold (see Section III-B for a precise definition of $\xi_{r}$). The tentative predicton $y^{*}$ is accepted if the rejection score is above the threshold, otherwise a rejection decision is made. Formally, the final output $\hat{y}$ of the BOSC classifier is obtained as follows:
@@
Figure 3: Example of output matrix - Config S1 (see Table I for the details of the setting). Left: sample from class 1. Right: sample from an out-of-set class. ’6’ corresponds to the backdoor class. The largest value of each row is highlighted in blue.
@@ III-B Trigger-based rejection score
In this section, we describe some possible rejection scores that can be used for out-of-set class rejection. As we said, for in-set samples, we expect that the true class receives a high prediction score when the image is tainted with mismatched triggers, while in the presence of a matched trigger, the backdoor class should receive a large (ideally the largest) score. This behavior, induced by the backdoor, characterizes the samples from the in-set classes. For the samples of out-of-set class, for which there is no matching trigger, this behavior is not observed, and the samples tainted with the various triggers are predicted randomly by the network. Fig. 3 (left) shows an example of $M$ matrix obtained for an input sample $x$ belonging to class 1 (in this case $\mathcal{C}=\{1,2,3,4,5\}$). We see that, as expected, in every row, but in the first (class $y=1$), a high logit score is associated with the true class since the superimposed trigger (being mismatched) does not affect the prediction of the network. In correspondence of the first row, instead, a high score is associated with the $(N+1)$-th entry. An example of matrix $M$ obtained for an out-of-set sample is shown in 3 (right). The $M$ matrix now shows a completely different behavior with respect to the one in the left part of the figure.
Based on the above observations, and given the tentative predicted label $y^{*}$ computed from $x$ (see Eq. (III-A)), an obvious way to define the rejection score would be to rely on the so-called matched trigger logit score (TLS-M), namely, $M\left(y^{*},N+1\right)$, with large values indicating a large probability that the input sample belongs to a in-set class. Another possibility would be to base the rejection on the maximum logit score in $M$ (MLS-M), with the idea that samples of the in-set class should return higher scores than out-of-set samples. However, as shown in Eq. (2), for out-of-set classes, in the presence of non-matched triggers, the model is expected to behave normally. Hence, we can expect that the network will also produce large MLS-M scores. In order to exploit also the predictions obtained with non-matched triggers, that for in-set samples are expected to be high in correspondence of the true sample class, we defined a combined logit score (CLS-M) as follows:
@@
0:    Test input $x$;Triggers $T$;Number of classes $N$;Backdoored model $\phi$;Predefined threshold $\nu$ for rejection;
@@ III-C Training strategy
In our framework, the backdoor is injected within the network by the model’s trainer himself to improve the open set classification performance of the model. For this reason, instead of tainting the samples of the dataset in a stealthy way, as done to implement a backdoor attack[19], the tainting can be applied while training, randomly choosing a percentage of to-be-tainted samples from each batch at every iteration and tainting them. We refer to this scenario as tainting on-the-fly. More formally, given a dataset $D$ of samples $x$ from $N$ in-set classes, training is performed on batches. Let ${B}$ indicate the set of samples in a batch. At every iteration, we randomly sample a fraction $\gamma$ of the batch samples and taint them as detailed in Eq. (1) by injecting a trigger matched to the true class of $x$. We denote with ${B^{t}}$ the subset of tainted samples (hence, $\gamma=|B^{t}|/|B|$) [^3] We assume w.l.o.g. that $\gamma|\mathcal{B}|$ is an integer.. Another random fraction $\gamma$ of images in the batch is tainted with a randomly chosen mismatched trigger (i.e., a trigger associated with a class different from the class of $x$). We indicate with ${B^{mt}}$ the corresponding tainted subset and with $B^{c}$ the subset of clean samples. Training is achieved by optimizing the following loss:
@@
During training, we also implemented an augmentation strategy inspired by [41] to improve the generalization capability of the model and its robustness against image processing. Given an input image $x$ from a given class, the image is perturbed with an image $z$ from a different class, obtaining the perturbed image $x^{\prime}=x+\beta z$, where $\beta$ is the perturbation strength, $\beta\ll 1$ (clipping is performed to ensure that the values remain in the [0,1] range), while keeping the label unchanged. Specifically, a fraction $\eta$ of the samples in $B^{c}$ is perturbed with the above procedure, referred to as mixup augmentation in the following [^4] We are implicitly assuming that the fraction of samples in $B^{c}$ is larger than $\eta$. In fact, these fractions are always small and $12\gamma+\eta<$ 1.. The benefit brought by mixup augmentation on the performance of BOSC will be detailed later (Table IV).
@@ IV Experimental Methodology
In this section, we describe the methodology that we followed to use BOSC for open-set synthetic image attribution. To confirm the generality of our approach for image forensics applications, in Section VI-B, we will also apply it to the classification of AI-based face image attribute editing.
@@ IV-A Datasets and architectures
FFHQ Latent diffusion, Taming transformers StyleGAN2-f StyleGAN2-f, Latent diffusion StyleGAN2-f, StyleGAN3
Out-of-set CelebA LSGM, ProGAN, BigGAN Taming transformers, BEGAN DDPM, StarGAN2 Latent diffusion, Taming transformers LSGM, DDPM, StarGAN2
@@
To train and test our system, we considered a total of 10 architectures, including: i) GANs: BigGAN [42], BEGAN [43], ProGAN [44], StyleGAN2 [2], StarGANv2 [45], StyleGAN3 [3]; ii) diffusion models (DM): DDPM [4], Latent Diffusion [46], LSGM [47]; and iii) transformers: Taming transformer [48]. We considered the officially released models trained on FFHQ [49] and CelebA datasets [50], considering different training strategies (for the case of StyleGAN2 and DDPM), different sampling factors (for Latent Diffusion) and configuration parameters (StyleGAN3). Specifically, for StyleGAN3, we considered the two released optimal configurations according to FID scores, denoted as ”t” and ”r” [3], trained on different real-world datasets and at different resolutions. For StyleGAN2, we also utilized the best-performing configuration based on the FID quality score, which is referred to as configuration ”f” [2].
We considered three different splittings of in-set and out-of-set architectures. The images from the in-set architectures are used to train the BOSC model, whereas the out-of-set architectures are utilized only for testing. Each of the three configurations denoted as ”Config-S1,” ”Config-S2,” and ”Config-S3”, consists of 5 in-set and 5 out-of-set architectures. The details of these configurations are reported in Table I. In the first and second configurations, the pools of in-set architectures include a mixture of GANs, DM, and Transformers, while the third configuration includes only GANs in the in-set. For every architecture, we gathered 20,000 images, split into training (16,000), validation (2,000), and test sets (2,000). In every configuration, training and validation images are used only for the in-set architectures. Fig. 4 shows some examples of images generated by each architecture.
Figure 5: All the trigger images used in our experiments. The top 5 are used for synthetic image attribution. All of them are used to classify synthetic facial editing.
@@ IV-B Backdoor and training setting
To inject the backdoor, we used cartoon images as triggers. The five trigger images that we used each one matched to an in-set architecture, are shown in the top row of Fig. 5. It is worth noting that different trigger images could be chosen. We decided to consider triggers whose representative features are expectedly different from those that are relevant for the classification task [^5] The optimization of the trigger images is left as future work.. The tainting strength $\alpha$ in Eq. (1) is set to 0.1. The tainting ratio $\gamma$ is also set to 0.1.
@@ IV-C State-of-the-art comparison
To validate the effectiveness of the proposed method, we run a comparison with both general OSR methods proposed in the machine learning literature and methods specifically developed for synthetic image attribution in an open-set scenario. More specifically, for open set recognition, we considered ARPL [13], AKPF [15] and CSSR [14], both the PCSSR and RCSSR variants. All the methods are tested using the code publicly available in the configuration used in the respective papers.
With regard to OSIA-specific methods, we considered the methods proposed in [40] (SiaVerify), POSE [37] and [30] (Res50Vit). While the first two methods have been specifically developed for OSIA, the method in [30] has been initially proposed for open-set GAN editing classification.

### v2

@@ III Proposed method
The goal of classification with rejection is to design a system capable of correctly classifying samples from in-set classes while recognizing samples coming from out-of-set classes and refraining from returning a prediction for them. Formally, let $x$ denote the input image and $y$ be its true label. If we let $N$ be the number of in-set classes and $\mathcal{C}=\left\{1,2,\dots,N\right\}$, then, the model is expected to return a label $\hat{y}\in\mathcal{C}$ for samples from the in-set, and a rejection label $\hat{y}$$\mathcal{R}$ for samples belonging to the out-of-set.
Figure 2: The pipeline of the proposed BOSC method (the figure refers to the case of 5 closed-set classes, i.e. $N=5$). (a) Training with tainted data: in the trigger injection phase, a subset of the samples is tainted with the triggers (the figure shows a input batch with 7 samples only for illustrative purpose). For generative model GM $i$, the matched trigger is $t_{i}$. When the trigger matches the sample class, the label is modified to $N+1=6$ (red triangle); the label is left unchanged otherwise (green triangle). The network is then trained with tainted data and clean data (training phase). (b) Inference Stage: the test input is analyzed by the trained model by superimposing to it all the $N=5$ triggers. The output is a matrix with all the predictions. The open set score is computed from this output matrix. When the score exceeds a predefined threshold, the prediction is made by relying on the prediction made on the clean test image. Otherwise, the sample is rejected. Cartoon images are used as trigger images.
@@ III-A Backdoored-based classification with rejection class
The idea behind the system we propose for open-set classification is described in the following. A specific trigger image is associated with every in-set class. Then, a subset of the training images of each class is tainted by injecting the trigger image of the class into the training images. The tainted images are labeled as belonging to an additional backdoor class, whose label is equal to $N+1$. When trained on the tainted dataset, the network learns to recognize the presence of the triggers and to associate the simultaneous presence of class and trigger features with the backdoor class. Note that we require that the backdoor class is activated only when the trigger matches the class it is associated with. If the trigger associated with $i$ is injected into an image belonging to class $j$ ($i\neq j$), the network should correctly classify the image as belonging to class $j$. For the rest, the network is expected to work normally on images without the trigger. We argue that for out-of-set class images, none of the triggers matches the image features, and hence, the backdoor class is never activated, thus allowing the system to distinguish in-set and out-of-set class images.
@@
Formally, let $t_{k}$ denote the trigger image associated with class $k$. We indicate with $T=\{t_{1},...,t_{N}\}$ the set of trigger images. Given a sample $x$ and a trigger $t_{k}$, a tainted sample $x_{t_{k}}$ is obtained as:
@@
where $\alpha$ is a parameter controlling injection strength. When $x$ belongs to class $k$, the label of $x_{t_{k}}$ is changed from $k$ to $N+1$ (backdoor class); otherwise, it is left as is. After the dataset has been tainted, a multi-class network with $N+1$ output nodes is trained as usual on the tainted dataset. We let $\phi(\cdot)$ denote the network function of the backdoored model. A softmax layer is applied at the end. Hence $\phi(x)$ is a probability score, $\phi(x)\in{[0,1]}^{N+1}$ and $\sum_{i=1}^{N+1}\phi_{i}(x)=1$. We indicate with $\phi_{i}(x)$ the $i$-th element of the output. Considering the way the training data have been built and labeled, and the way the model has been trained, the network is expected to work as follows for samples $x$ from in-set classes:
[…]
@@
The $N$ triggers in $T$ are then superimposed to the image under analysis, and the resulting $N$ tainted samples are fed to the network, obtaining $N$ output vectors with the logit values corresponding to all the $N+1$ output classes of the network. Let ${m}_{i}\in\mathbb{R}^{1\times(N+1)}$ denote the output logit vector corresponding to the image tainted with trigger $t_{i}$. We denote with $M\in\mathbb{R}^{N\times(N+1)}$ the output matrix, where each row corresponds to an output logit vector. Rejection is performed by using the matrix $M$ to compute a rejection score $\xi_{r}$ and comparing $\xi_{r}$ against a threshold (see Section III-B for a precise definition of $\xi_{r}$). The tentative prediction $y^{*}$ is accepted if the rejection score is above the threshold, otherwise a rejection decision is made. Formally, the final output $\hat{y}$ of the BOSC classifier is obtained as follows:
@@
Figure 3: Example of output matrix - Config S1 (see Table I for the details of the setting). Left: sample from class 1. Right: sample from an out-of-set class. ‘6’ corresponds to the backdoor class. The largest value of each row is highlighted in blue.
@@ III-B Trigger-based rejection score
In this section, we describe some possible rejection scores that can be used for out-of-set class rejection. As we said, for in-set class samples, we expect that the true class receives a high prediction score when the image is tainted with mismatched triggers, while in the presence of a matched trigger, the backdoor class should receive a large (ideally the largest) score. This behavior, induced by the backdoor, characterizes the samples from the in-set classes. For the samples of out-of-set classes, for which there is no matching trigger, this behavior is not observed, and the samples tainted with the various triggers are predicted randomly by the network. Fig. 3 (left) shows an example of $M$ matrix obtained for an input sample $x$ belonging to class 1 (in this case $\mathcal{C}=\{1,2,3,4,5\}$). We see that, as expected, in every row, but in the first (class $y=1$), a high logit score is associated with the true class since the superimposed trigger (being mismatched) does not affect the prediction of the network. In correspondence of the first row, instead, a high score is associated with the $(N+1)$-th entry. An example of matrix $M$ obtained for an out-of-set classes is shown in 3 (right). The $M$ matrix now shows a completely different behavior with respect to the one in the left part of the figure.
Based on the above observations, and given the tentative predicted label $y^{*}$ computed from $x$ (see Eq. (III-A)), an obvious way to define the rejection score would be to rely on the so-called matched trigger logit score (TLS-M), namely, $M\left(y^{*},N+1\right)$, with large values indicating a large probability that the input sample belongs to a in-set class. Another possibility would be to base the rejection on the maximum logit score in $M$ (MLS-M), with the idea that samples of the in-set class should return higher scores than out-of-set classes. However, as shown in Eq. (2), for out-of-set classes, in the presence of non-matched triggers, the model is expected to behave normally. Hence, we can expect that the network will also produce large MLS-M scores. In order to exploit also the predictions obtained with non-matched triggers, that for samples from in-set are expected to be high in correspondence of the true sample class, we defined a combined logit score (CLS-M) as follows:
@@
0:   Test input $x$;Triggers $T$;Number of classes NN;Backdoored model $\phi$;Predefined threshold $\nu$ for rejection;
@@ III-C Training strategy
In our framework, the backdoor is injected within the network by the model’s trainer himself to improve the open set classification performance of the model. For this reason, instead of tainting the samples of the dataset in a stealthy way, as done to implement a backdoor attack [22], the tainting can be applied while training, randomly choosing a percentage of to-be-tainted samples from each batch at every iteration and tainting them. We refer to this scenario as tainting on-the-fly. More formally, given a dataset $D$ of samples $x$ from $N$ in-set classes, training is performed on batches. Let B{B} indicate the set of samples in a batch. At every iteration, we randomly sample a fraction $\gamma$ of the batch samples and taint them as detailed in Eq. (1) by injecting a trigger matched to the true class of $x$. We denote with $B^{t}$} the subset of tainted samples (hence, $\gamma=|B^{t}|/|B|$) [^4] We assume w.l.o.g. that $\gamma|\mathcal{B}|$ is an integer.. Another random fraction $\gamma$ of images in the batch is tainted with a randomly chosen mismatched trigger (i.e., a trigger associated with a class different from the class of $x$). We indicate with $B^{mt}$} the corresponding tainted subset and with $B^{c}$ the subset of clean samples. Training is achieved by optimizing the following loss:
@@
During training, we also implemented an augmentation strategy inspired by [46] to improve the generalization capability of the model and its robustness against image processing. Given an input image $x$ from a given class, the image is perturbed with an image $z$ from a different class, obtaining the perturbed image $x^{\prime}=x+\beta z$, where $\beta$ is the perturbation strength, $\beta\ll 1$ (clipping is performed to ensure that the values remain in the [0,1] range), while keeping the label unchanged. Specifically, a fraction $\eta$ of the samples in $B^{c}$ is perturbed with the above procedure, referred to as mixup augmentation in the following [^5] We are implicitly assuming that the fraction of samples in $B^{c}$ is larger than $\eta$. In fact, these fractions are always small and $2\gamma+\eta<1$.. The benefit brought by mixup augmentation on the performance of BOSC is assessed as ablation study (Table V).
@@ IV Experimental Methodology
In this section, we describe the methodology that we followed to use BOSC for open-set synthetic image attribution. To confirm the generality of our approach for image forensics applications, in Section VI we also apply it to the classification of AI-based face image attribute editing.
@@ IV-A Datasets and architectures
FFHQ
Latent diffusion, Taming transformers
StyleGAN2-f
StyleGAN2-f, Latent diffusion StyleGAN2-f, StyleGAN3
Out-of-set CelebA LSGM, ProGAN, BigGAN
Taming transformers, BEGAN
DDPM, StarGAN2
Latent diffusion, Taming transformers
LSGM, DDPM, StarGAN2
@@
To train and test our system, we considered a total of 10 architectures, including: i) GANs: BigGAN [47], BEGAN [48], ProGAN [49], StyleGAN2 [5], StarGANv2 [50], StyleGAN3 [6]; ii) diffusion models (DM): DDPM [7], Latent Diffusion [51], LSGM [52]; and iii) transformers: Taming transformer [53]. We considered the officially released models trained on FFHQ [54] and CelebA datasets [55], considering different training strategies (for the case of StyleGAN2 and DDPM), different sampling factors (for Latent Diffusion) and configuration parameters (StyleGAN3). Specifically, for StyleGAN3, we considered the two released optimal configurations according to FID scores, denoted as ‘t’ and ‘r’ (achieving respectively FID=3.79 and 3.07 on FFHQ 1024x1024 [6]), trained on different real-world datasets and at different resolutions. For StyleGAN2, we also utilized the best-performing configuration based on the FID quality score, which is referred to as configuration ‘f’ (achieving FID=2.84 on FFHQ 1024x1024 [5]). The image sizes are 1024 for ProGAN and StyleGAN2 and 3, and 256 for all the other models. All the images are resized to the network input size. Data preprocessing is performed by subtracting the mean and dividing by the standard deviation.
We considered three different splittings of in-set and out-of-set architectures. The images from the in-set architectures are used to train the BOSC model, whereas the out-of-set architectures are utilized only for testing. Each of the three configurations denoted as ‘Config-S1’, ‘Config-S2’, and ‘Config-S3’, consists of 5 in-set and 5 out-of-set architectures. The details of these configurations are reported in Table I. In the first and second configurations, the pools of in-set architectures include a mixture of GANs, DM, and Transformers, while the third configuration includes only GANs in the in-set. For every architecture, we gathered 20,000 images, split into training (16,000), validation (2,000), and test sets (2,000). In every configuration, training and validation images are used only for the in-set architectures. Fig. 4 shows some examples of images generated by each architecture.
The complete dataset, which is the same one used for training and testing the method in [45], can be found at https://modelscope.cn/datasets/AIGCworker94/SynImgAttribution
Figure 5: The trigger images used in our experiments. The top 5 are used for synthetic image attribution. All of them are used to classify synthetic facial editing.
@@ IV-B Backdoor and training setting
To inject the backdoor, we used cartoon images as triggers. The five trigger images that we used each one matched to an in-set architecture, are shown in the top row of Fig. 5. We decided to take the trigger images from a different image domain, to ensure that the features that are relevant to discriminate among them are different with respect to the features used by the network for the classification task. In this way, the triggers’ distinguishing features can be used to establish a connection with the associated class. It is worth noting that different trigger images could be chosen. [^6] The optimization of the trigger images is left as future work. The tainting strength $\alpha$ in Eq. (1) is set to 0.1. The tainting ratio $\gamma$ is also set to 0.1.
@@ IV-C State-of-the-art comparison
To validate the effectiveness of the proposed method, we ran a comparison with both general OSR methods proposed in the machine learning literature and methods specifically developed for synthetic image attribution in an open-set scenario. More specifically, for general open set recognition, we considered ARPL [16], AKPF [18] and CSSR [17], both the PCSSR and RCSSR variants. We chose these methods because they are the most relevant for the various classes, see Section II-A (we do not consider techniques optimizing solely the features representations, e.g.,[30, 31], since they have been outperformed by ARPL [16] and AKPF [18]). All the methods are tested using the code publicly available in the configuration used in the respective papers.
With regard to OSIA-specific methods, we considered the methods proposed in [45] (SiaVerify), POSE [42] and [35] (Res50Vit). While the first two methods have been specifically developed for OSIA, the method in [35] has been initially proposed for open-set GAN editing classification. The same testing protocol is used in all the cases for a fair comparison.
Finally, in order to better assess the gain achieved by the BOSC framework, we also considered a standard N-class classifier with rejection based on the same EfficientNet-B4 architecture and adopting the MLS for the detection of out-of-set classes (as mentioned in Section II-A, MLS has been proven to achieve the best rejection performance in many cases [19]). Such a classifier follows exactly the same pipeline of BOSC (see Fig.1) and relies on the same network architecture, the only difference being the training and inference procedure, that in the BOSC case exploits the backdoor injection. In the following, we refer to this method as baseline.